What amount of each mixture, one 95% alcohol the other 15% alcohol must be used to make 10 liters of a mixture which is 45% alcohol?
Answer by stanbon(75887)

What amount of each mixture, one 95% alcohol the other 15% alcohol must be used to make 10 liters of a mixture which is 45% alcohol?
--------------------------
Equation:
alcohol + alcohol = alcohol
0.95x + 0.15(10-x) = 0.45*10
Multiply thru by 100 to get:
95x + 150 - 15x = 450
80x = 300
x = 3.75 liters (amt. of 95% solution)
10-x = 6.25 liters (amt. of 15% solution)
================================================
Cheers,
stan H.